A vegetable gardener has a tilled a piece of soil that measures 6 yards by 12 yards. She wants to put a border of marigolds of the same width within the four sides and leave an area of 40 square yards for vegetables. How wide can the border be?
Answer by macston(5194)

.
(6yds-x)(12yds-x)=40yds^2
72yds-6xyds-12xyds+x^2=40yds^2
x^2-18xyds+72yds^2=40yds^2
x^2-18xyds+32yds=0
(x-2yds)(x-16yds)=0
x-2yds=0 OR x-16yds=0
x=2yds OR x=16yds x=16yds results in negative measurements, so
x=2 yards ANSWER: The border can be 2 yards wide.